459 laser 30

A laser beam traces a path through a prism, as shown. Assume the room is filled with air and the prism is made of some unknown clear material. Find the index of refraction of the prism.
Answer #1

Mair=1 459 8 r n - L$+. 45° 90° = Lo 2° = 90 - 45° 45° 18+ 24 = 90° = 28 -90-60=90-45° = 45° o =30° Letco 90° Le = 90° -30° 6